OpenStudy (anonymous):

George folds his origami paper along the dotted lines shown below so that the triangular flaps overlap exactly on the square. If the dimensions of the origami paper are decreased by One-half, what will be the new area of the paper? Answers A. 98 square inches <----My choice B. 196 square inches C. 49 square inches D. 147 square inches

OpenStudy (anonymous):

whats below if its a pic i need to see it

OpenStudy (anonymous):

I was uploading it...sorry. It's up now.

OpenStudy (anonymous):

Its 9.8 (approximately) are there supposed to be decimal points in those answers? if so you're right. Good job!

OpenStudy (anonymous):

Thanks, I might have more questions in a few minutes would you be willing to help?

OpenStudy (anonymous):

Bloodrain is wrong, the answer is 49

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

2x^2 = (7sqrt2)^2 (pythagorean theorem) 2x^2=49*2 x^2=49 x=7 new area = x^2 = 49

OpenStudy (anonymous):

Oh, got it. Thanks

OpenStudy (anonymous):

Wait which area do you need?

OpenStudy (anonymous):

The second one.

OpenStudy (anonymous):

oh nvm then it's just (7sqrt2)^2=98

OpenStudy (anonymous):

So @bloodrain and I are right?

OpenStudy (anonymous):

Yeah, i misunderstood the question

OpenStudy (anonymous):

But then there is no point of even providing the info about origami, it should just say to find the area of a square with that side length. The intro was unnecessary

OpenStudy (anonymous):

yeah its because its a story problem and they are supposed to be confusing